OpenStudy (anonymous):

Solve the formula for the variable y. z = 4x + 2y @E.ali

OpenStudy (anonymous):

2y= z - 4x y= 1/2 (z - 4x)

OpenStudy (anonymous):

We can say : z-4x=2y z/2-2x=y and ... can you say the answers ?:)

OpenStudy (anonymous):

A. y = z/2 - 4x B. y = z/2 - 2x C. y = z/2 + 2x D. y = z - 2x

OpenStudy (anonymous):

y= 1/2 (z - 4x) Is : y= 1/2 (z - 4x)

OpenStudy (anonymous):

I said ! It s B .!

OpenStudy (anonymous):

thankyou!! you just helped me complete my test lets see what got!!!

OpenStudy (anonymous):

K !

OpenStudy (anonymous):

i failed it!

OpenStudy (anonymous):

i got 8/12 wich is 66.7%

OpenStudy (anonymous):

i got to bounus question if i get these right thn i pass